OpenStudy (anonymous):

find the indicated sum for the arithmetic series S18 for 3.2+2.9+2.6+2.3+...

OpenStudy (goldeverything):

do you know how to solve this equation

OpenStudy (anonymous):

yeah i already found the common difference which is 0.3

OpenStudy (goldeverything):

so what do you do after that

OpenStudy (anonymous):

i put the equation: a18=3.2+(18-1)(0.3) =8.3

OpenStudy (goldeverything):

correct

OpenStudy (anonymous):

then i got the equation Sn=n[ a1+a2/2] plugged in the numbers S18 = 18 [3.2+8.3/2] =18(5.75) =103.5

OpenStudy (goldeverything):

yes that is the awnser good job

OpenStudy (anonymous):

really ? cool, thanks man(:

OpenStudy (goldeverything):

its no problem please message me if you have any further questions